OpenStudy (anonymous):

Convert to a decimal: 1 7/8 1.825 1.9 1.875 1.85

OpenStudy (godlovesme):



OpenStudy (godlovesme):

multiply 1 and 8 = 8 then add 8+7 = 15 15/8 divide 15 by 8 @sunnyo

OpenStudy (anonymous):

1.875

OpenStudy (anonymous):

1 7/8 = 1.875 Calculation: 1 7/8 = 1 + 7/8 = 1/1 + 7/8 = (1/1 * 8/8) + 7/8 = 8/8 + 7/8 = 15/8 = 15 ÷ 8 = 1.875 Alternate Calculation: 1 7/8 = 1 + 7/8 = 1 + (7 ÷ 8) = 1 + 0.875 = 1.875 @Sunnyo

OpenStudy (anonymous):

1.875

OpenStudy (anonymous):

Thank you all!

OpenStudy (godlovesme):

yw :)

OpenStudy (anonymous):

NP